# G-Loc: Tightly-coupled Graph Localization with Prior Topo-metric Information

### v1

@@ IV-A LiDAR-inertial Odometry (LIO)
where ${}^{i}\tilde{\boldsymbol{T}}^{\text{LIO}}_{j}$ is the output from the previous subsection, ${}^{i}\tilde{\boldsymbol{T}}_{k}^{\mathcal{M}}$ are the registration results between the current scan and the prior map, and $\boldsymbol{T}^{\prime}_{k}\in\mathcal{G}$. Note that correspondences with the prior map are only evaluated for a subset of the reference frames, i.e., $k\in\mathcal{R}_{i}\subset\mathcal{R}$. Finally, $\tilde{\boldsymbol{p}}_{i}={}^{\text{UTM}}\tilde{\boldsymbol{p}}_{i}\texttt{-}{}^{\text{UTM}}\tilde{\boldsymbol{p}}_{m}$, with ${}^{\text{UTM}}\tilde{\boldsymbol{p}}_{i}$ the actual GNSS measurement and ${}^{\text{UTM}}\tilde{\boldsymbol{p}}_{m}$ the origin of the prior model (i.e., the GNSS estimation for the first vertex). Note that this origin will be associated with a UTM (Universal Transverse Mercator) coordinate. Relativizing GNSS allows to merge it with cloud-to-map matching and LIO constraints.
Dataset Length $\Delta$ Elevation LiDAR IMU GNSS
EU long-term (ref. & eval.) $\sim 5\phantom{.0}$ km $\sim 50$ m 2x Velodyne HDL-32E XSens MTi-28A53G25 Magellan ProFlex 500
UZ Campus (ref.) $\sim 2.7$ km $\sim\phantom{0}1$ m Ouster OS1-128 Ardusimple RTK2B
UZ Campus (eval.) $\sim 1.5$ km $\sim\phantom{0}1$ m Livox HAP
Bus $\sim 8\phantom{.0}$ km $\sim 70$ m 2x RS Helios 32-5515 XSens MTI-680-G
@@
On the one hand, LIO relative measurements are used to connect sequential vertices of the sliding graph with information depending on the alignment result (defined as the mean error between all matched points, weighted by the match percentage). In addition, it also provides an initial guess for the cloud-to-map registration. On the other hand, registrations between the current scan and the reference submaps are introduced as binary edges between vertices in the reference and sliding graphs, with information also depending on the alignment result. Also, the transformation with the highest information is used to compute the initial estimate for the optimization process (also handled by g2o). The previous constraints connect the two graphs and form the active graph (blue elements in Figure 1). Should the sliding and reference graphs become disconnected, either due to alignment failure or the lack of nearby reference vertices, the algorithm will rely on LIO and GNSS until re-localization occurs. Finally, GNSS measurements are used as prior unary edges (blue arrows in Figure 1). These constraints are only incorporated if they are synchronized with the online cloud and if their uncertainty is below a certain threshold, thus rejecting noisy measurements.
IV-C Dynamic Loading
A dynamic loader is used to relax the high memory requirements of loading large point cloud maps. This module allows loading only nearby parts of the reference model, $\mathcal{M}$, as the robot moves; this is especially useful for large maps where maintaining (and visualizing) all submaps at high resolution is intractable. Since both the reference and sliding graphs are geo-referenced, it is possible to load the submaps near the current robot localization estimate in a sliding window fashion. To do this, we query the nearby vertices in the reference graph (i.e., within a specified Euclidean distance) and keep only their associated submaps in memory. In Figure 4, only the submaps that are close to the active graph (shown in blue) are loaded into memory and thus visualized in a large environment.
V Experimental Results
((a))
((b))
((c))
((d))
Figure 5: Evaluation results in the EULT (a)-(b) and UZ Campus (c)-(d) datasets. (a), (c): geo-referenced SLAM trajectories in the reference sequences (orange) and raw GNSS data in the evaluation sequences (green). (b), (d): Trajectories estimated by G-Loc in the evaluation sequences; the color map indicates 2D ATE (darker is smaller).
In this section, we present a thorough evaluation of G-Loc in different applications and under different sensor configurations, demonstrating the versatility of our approach. We use two real urban datasets: EU long-term dataset [34] and our own dataset collected on the campus of the University of Zaragoza (UZ). In addition, we present the results of using G-Loc in a real-world automotive application, in the context of an autonomous driving bus. The dataset experiments were performed on a laptop equipped with an Intel Core i7 CPU (8 cores) and an Nvidia GeForce RTX 3080 GPU, while the bus was equipped with an Intel Core i9 CPU (24 cores) and a Nvidia A2000 GPU (note that many other processes for autonomous driving were running concurrently on this hardware). Unless otherwise noted, [31] was used for registration. Table I contains a short summary of the experiments and the sensors installed in the mobile platforms. The results presented in this section were obtained using evo [^3] https://github.com/MichaelGrupp/evo.
@@ V-A Dataset Evaluation
The EU long-term dataset contains several sequences recorded with a car in a downtown. They all follow a similar trajectory under different conditions (e.g., traffic, weather); this is particularly interesting for testing localization, since we need different but overlapping reference and evaluation trajectories. The length of each sequence is about $5$ km with a duration of $16$ minutes. The sensor configuration is detailed in Table I.
@@
To evaluate the performance of G-Loc on this dataset, we first generated a geo-referenced model of the environment using one of the available sequences, which will be referred to as reference sequence [^4] Reference sequence id: 2018-07-19 (Thu, sunny). Evaluation sequences ids: 2018-07-16 (Mon, sunny) and 2018-07-17 (Tue, sunny).. We used a SLAM system that inherits the architecture shown in Figure 2 but replaces the localization module with a mapping module that features robust loop closing and global optimization. This way, we obtained the required prior model with topological (a pose-graph) and metric (point cloud submaps) information. To evaluate the performance in map-less regions, we removed an entire road from the prior model (approximately $0.4$ km); the SLAM-estimated trajectory is shown in orange in Figure 5(a). The built model was fed into G-Loc while running two different evaluation sequences4. Figure 5(a) also shows the raw GNSS data (green) for one of the evaluation sequences to facilitate its comparison with the reference trajectory; the middle right part contains the cropped region, where there is GNSS signal but no prior information. Whenever the robot traverses a mapped region, the prior model is exploited to improve localization and correct for potential drift in LIO. In unmapped regions, the sliding graph disconnects from the reference graph, accumulating larger errors that are eventually corrected after re-localization. Figure 5(b) contains the 2D Absolute Trajectory Error (ATE) for translation (in meters) along the trajectory for one of the sequences555 Due to the lack of ground-truth in this dataset and the noise levels of GNSS in certain regions, errors were calculated w.r.t. SLAM. This algorithm fuses GNSS with LIO to mitigate the effects of noisy measurements and GNSS-denied areas (see zoomed region of Figure 5(a)). The error between GNSS and SLAM is negligible when the former is reliable., clearly showing that the highest errors occur on the unmapped road and that even under these conditions they are within reasonable limits. Table II contains the numerical results (mean and standard deviation) for this experiment, namely 3D ATE for translation and rotation, and lateral and longitudinal error. This table also reports the results of using a complete prior map (i.e., without cropping a road); in this case, the errors were similar but show a smaller deviation. We evaluated the use of parallelized NDT, which runs entirely on the CPU and is of particular interest for resource-constrained platforms. In this case, localization errors were slightly higher than with GPU acceleration.
Finally, to benchmark G-Loc, we conducted the same experiments using Autoware localization module [8], which represents a popular state-of-the-art localization system for autonomous vehicles. It is based on an extended Kalman filter that integrates velocity measurements with cloud-to-map alignments and a GNSS regularization term. For a fair comparison, we fed velocity measurements from our LIO module and used parallelized NDT for registration. Autoware failed to localize in the mapless region, leading to an unrecoverable state. Using the complete prior model, the translation errors were comparable to G-Loc but the rotational error was nearly doubled (see Table II). In addition, Autoware required using an Intel Core i9 CPU with 24 cores to work properly.
@@
EULT (cropped) G-Loc (GPU) 0.209 (0.395) 0.087 (0.280) 0.048 (0.228) 0.423 (0.445) 25.13 (7.82)
G-Loc (CPU) 0.240 (0.300) 0.086 (0.304) 0.021 (0.208) 1.008 (1.160) 38.01 (14.86)
@@
EULT (complete) G-Loc (GPU) 0.194 (0.201) 0.101 (0.240) 0.014 (0.197) 0.440 (0.576) 26.47 (4.31)
G-Loc (CPU) 0.218 (0.196) 0.119 (0.245) 0.009 (0.203) 0.953 (0.836) 40.13 (12.99)
@@
TABLE II: Localization errors (mean and std. dev.) and time consumed per frame for G-Loc (GPU and CPU) and Autoware in the evaluated datasets. (--) indicates an unrecoverable failure in localization. The best results for each dataset are shown in bold.
The UZ Campus dataset was captured using a sensor-equipped car. We recorded two sequences, which partially overlap, but with two different LiDAR sensors. See Table I.
To evaluate this dataset, we followed a similar procedure as before. First, we built the prior model with a reference sequence using a high-performance sensor with $360\degree$ Field of View (FoV) and $128$ planes. Then, we evaluated G-Loc with an evaluation sequence, exploiting the prior topo-metric knowledge and using a LiDAR with $120\degree$ FoV and a completely different sensing pattern. Figure 5(c) displays the trajectory of the reference sequence (orange) and the raw GNSS data for the evaluation one (green). Both trajectories overlap for most of the route, and when they do not, they are close enough so that the prior map can be used (unlike in EULT). ATE error along the trajectory is shown in Figure 5(d) and numerical results are reported in the bottom of Table II. As in EULT, the use of the GPU gave the best results, except for the lateral error (however, these lower mean values are accompanied by higher variance). This experiment demonstrates the adaptability of our method to the use of different sensors for building the prior map and localization. The submaps played a fundamental role here, mitigating the effects of the different sensing patterns and feeding the registration algorithms with sufficient overlapping regions. In contrast, Autoware accumulated much higher errors and had difficulty finding correspondences in many regions (which explains the large deviations), eventually leading to an unrecoverable state. The results reported in Table II for this experiment (marked with ) are only before this divergence.
@@ V-B On the Time Consumption
While the above section aimed to demonstrate the robustness and precision of our approach, we now seek to demonstrate its real-time performance. The processing times per frame consumed by each algorithm in the conducted experiments are listed in Table II. G-Loc required, on average, $25$ ms in EULT and $34$ ms UZ Campus —this is much faster than the typical LiDAR frame rate ($10$ Hz) and therefore meets real-time requirements. In both cases, LIO accounts for $60\%$ of that time, and the localization module for the rest (visualization and dynamic load are excluded as they operate in the background). The fact that two different sensors were used in the second dataset resulted in longer registration times. G-Loc without GPU took considerably longer, but the frame rate was generally not exceeded. We also report the results for Autoware (marked with ), which are similar to G-Loc without GPU. Note, however, that Autoware used a more powerful CPU and had 3 times the number of cores available to parallelize computations, so the direct comparison is unfair.
Finally, we compared the use of G-Loc to naively running a full SLAM system again to reveal the time benefits of exploiting the known prior topo-metric map instead of using a new map. We ran this experiment on the EULT dataset because it is the largest and therefore where a new map would be more costly to generate and maintain. At first glance, the time required to process each frame might seem similar to that of G-Loc (the LIO update), but the graph optimization and loop closing threads took $90.4\pm 101.4$ ms (this scales with the graph size, unlike in G-Loc that optimizes a bounded-sized sliding graph) and $100.2\pm 0.4$ ms, respectively. These demanding processes usually slowed down the entire pipeline and caused interruptions in the localization estimates.
@@ V-C Real-world Deployment
G-Loc has been deployed on an autonomous bus (see Figure 6(a)) on an urban bus line in the city of Zaragoza (Aragón, Spain) as part of the pioneering Spanish project on urban driving, DIGIZITY. The sensors installed on the vehicle and the route specifics appear in Table I. Our system requires coordination with other navigation systems to provide a robust and reliable localization, ensuring safe navigation even in areas with poor GNSS coverage due to large buildings; accurate localization is critical as the bus must autonomously stop at designated bus stops and obey traffic lights. The prior model of the entire route ($\sim 8$ km) is shown in Figure 6(b). In addition, localization starts as soon as the bus is turned on. This means that the localization has to work without a prior model from the departure point until the route is reached ($\sim 2$ km). The system localizes in the unmapped region using LIO and GNSS in a sliding-graph optimization fashion until re-localization with the prior model occurs. Figure 6(c) depicts the trajectory estimated by G-Loc for a single trip, and Figure 6(d) shows the improvement of these estimates (blue) over the raw GNSS data (green) in one of the many low coverage areas. During the deployment, the bus operated seamlessly for about $100$ hours over a span of $20$ days, covering about $800$ km. Notably, there were no localization failures and the errors remained within bounds comparable to previous results (i.e., at the cm level), further showcasing the robustness and accuracy of G-Loc.
((a))
((b))
((c))
((d))

### v2

@@ IV-A LiDAR-inertial Odometry (LIO)
where ${}^{i}\tilde{\boldsymbol{T}}^{\text{LIO}}_{j}$ is the output from the previous subsection, ${}^{i}\tilde{\boldsymbol{T}}_{k}^{\mathcal{M}}$ are the registration results between the current scan and the prior map, and $\boldsymbol{T}^{\prime}_{k}\in\mathcal{G}$. Note that correspondences with the prior map are only evaluated for a subset of the reference vertices, i.e., $k\in\mathcal{R}_{i}\subset\mathcal{R}$. Finally, $\tilde{\boldsymbol{p}}_{i}={}^{\text{UTM}}\tilde{\boldsymbol{p}}_{i}\texttt{-}{}^{\text{UTM}}\tilde{\boldsymbol{p}}_{m}$, with ${}^{\text{UTM}}\tilde{\boldsymbol{p}}_{i}$ the actual GNSS measurement and ${}^{\text{UTM}}\tilde{\boldsymbol{p}}_{m}$ the origin of the prior model (this origin will be associated with a Universal Transverse Mercator coordinate). Relativizing GNSS allows to merge it with cloud-to-map matching and LIO constraints.
Figure 5: Visualization of dynamic loading. The prior graph is shown in orange, while the active graph is depicted in blue. For efficiency, only a subset of the geometric model is loaded, specifically the submaps corresponding to vertices within the robot’s neighborhood.
On the one hand, LIO is used to connect sequential vertices of the sliding graph with the covariance depending on the alignment result (defined as the mean error between all matched points, weighted by the match percentage). In addition, it provides an initial guess for the cloud-to-map registration. On the other hand, registrations between the current scan and the reference submaps are introduced as binary edges between vertices in the reference and sliding graphs, with information also depending on the alignment result. Also, the transformation with the smallest covariance is used to compute the initial estimate for the optimization process (handled by g2o). The previous constraints connect the two graphs and form the active graph (blue elements in Figure 4). Should the sliding and reference graphs become disconnected, either due to alignment failure or the lack of nearby reference vertices, the algorithm will rely on LIO and GNSS until re-localization occurs. Finally, GNSS measurements are used as prior unary edges when they are synchronized with the online cloud and if their uncertainty is below a certain threshold.
IV-C Dynamic Loading
[…]
V Experimental Results
Dataset Length (per sequence) $\Delta$ Elevation LiDAR IMU GNSS
EULT (ref. & eval.) $5\phantom{.0}$ km $50$ m 2x Velodyne HDL-32E XSens MTi-28A53G25 Magellan ProFlex 500
NCE (ref. & eval.) $0.3$ km $\phantom{0}0$ m Ouster OS0-128 --
CUZ (ref.) $2.7$ km $\phantom{0}1$ m Ouster OS1-128 Ardusimple RTK2B
CUZ (eval.) $1.5$ km $\phantom{0}1$ m Livox HAP
Bus $8\phantom{.0}$ km $70$ m 2x RS Helios 32-5515 XSens MTI-680-G
@@ V Experimental Results
Figure 6: Evaluation results in EULT (a)-(b) and CUZ (c)-(d) datasets. (a), (c): georeferenced SLAM trajectories in the reference sequences (orange) and raw GNSS data in the evaluation sequences (green). (b), (d): Trajectories estimated by G-Loc in the evaluation sequences; the color map indicates 2D ATE (in meters, darker is smaller).
[…]
@@ V-A Dataset Evaluation
The EULT dataset contains several sequences recorded with a car in a downtown. They all follow a similar trajectory under different conditions (e.g., traffic, weather); this is particularly interesting for testing map-based localization, since we need different but overlapping reference and evaluation trajectories. Sequence and sensor details are shown in Table I.
@@
To evaluate the performance of G-Loc, we first generated a georeferenced model of the environment using one of the available sequences, which we will refer to as the reference sequence [^2] Ref. seq. ID: 2018-07-19. Eval. seq. ID: 2018-07-16 and 2018-07-17.. To obtain the required prior model with topological (a pose-graph) and metric (point cloud submaps) information, we used [29], a SLAM system with a similar architecture to G-Loc but replacing the localization module with a mapping module for robust loop closing and global optimization. To evaluate the performance in mapless regions, we removed an entire road from the prior model ($0.4$ km); the SLAM-estimated trajectory is shown in orange in Figure 6a. The built model was fed into G-Loc while running two different evaluation sequences2. Figure 6a also depicts the raw GNSS data (green) for one of them to facilitate comparison with the reference trajectory. The middle right section contains the cropped region, where there is GNSS signal but no prior information. When the robot traverses a mapped region, the prior model is exploited to improve localization and correct for potential drift in LIO. Conversely, in unmapped regions, the sliding graph disconnects from the reference graph, accumulating a larger error that is eventually corrected after re-localization. Global measurements in this large mapless region kept the absolute error bounded, enabling the system to re-localize with the map as it traversed previously mapped areas. Figure 6b contains the 2D Absolute Trajectory Error (ATE) for translation along the trajectory for one of the sequences333 Due to the lack of ground-truth in this dataset and the noise levels of GNSS in certain regions, errors were calculated w.r.t. SLAM. This algorithm fuses GNSS with LIO to mitigate the effects of noisy measurements and GNSS-denied areas (see zoomed region of Figure 6a). The error between GNSS and SLAM is negligible when the former is reliable., showing that the highest errors occur on the unmapped road and that even under these conditions they are within reasonable limits (see colormap). Table II contains the numerical results (mean and standard deviation) for the cropped sequences, namely Root Mean Squared Error (RMSE) of 3D ATE for translation and rotation, and lateral and longitudinal error. Furthermore, we report the results of using a complete prior map (i.e., uncropped); in this case, the errors were similar but exhibited a smaller deviation. We also evaluated the use of NDTOmp, which is of particular interest for resource-constrained platforms. In this case, the overhead slightly increased the localization errors.
[…]
@@
EULT (cropped) G-Loc (FastGICP) 0.209 (0.395) 0.087 (0.280) 0.048 (0.228) 0.423 (0.445) 25.13 (7.82)
G-Loc (NDTOmp) 0.240 (0.300) 0.086 (0.304) 0.021 (0.208) 1.008 (1.160) 38.01 (14.86)
@@
EULT (complete) G-Loc (FastGICP) 0.194 (0.201) 0.101 (0.240) 0.014 (0.197) 0.440 (0.576) 26.47 (4.31)
G-Loc (NDTOmp) 0.218 (0.196) 0.119 (0.245) 0.009 (0.203) 0.953 (0.836) 40.13 (12.99)
@@
TABLE II: Localization errors (mean and std. dev.) and time consumed per frame for G-Loc (FastGICP and NDTOmp) and Autoware in EULT and CUZ datasets. Best results are bold, and $\times$ indicates an unrecoverable failure in localization.
The CUZ dataset was captured using a sensor-equipped car. We recorded two sequences, which partially overlap, but with two different LiDAR sensors. See Table I. To evaluate this dataset, we followed a similar procedure as in EULT. First, we built the prior model with a reference sequence using a high-performance sensor with $360\degree$ Field of View (FoV) and $128$ planes. Then, we evaluated G-Loc with an evaluation sequence, exploiting the prior topo-metric knowledge and using a LiDAR with $120\degree$ FoV and a completely different sensing pattern. Figure 6c displays the trajectory of the sequence (orange) and the raw GNSS data for the evaluation one (green). Both trajectories overlap for most of the route, and when they do not, they are close enough so that the prior map can be used (unlike in EULT). ATE error along the trajectory is shown in Figure 6d and numerical results are reported in the bottom of Table II. As in EULT, the use of FastGICP gave the best results, except for the lateral error (however, these lower mean values are accompanied by higher variance). This experiment demonstrates the adaptability of our method to the use of different sensors for building the prior map and localization. The submaps played a fundamental role here, mitigating the effects of the different sensing patterns and feeding the registration algorithms with sufficient overlapping regions. In contrast, Autoware accumulated much higher errors and had difficulty finding correspondences in many regions (which explains the large deviations), eventually leading to an unrecoverable state. The results reported in Table II for this experiment (marked with ) are only before this divergence.
The NCE dataset contains several sequences recorded using a handheld device in Oxford. This dataset allows us to test G-Loc under different conditions than the previous ones. It includes aggressive motion, where IMU integration is crucial for undistorting the point cloud and predicting the motion between successive LiDAR frames. We have used Quad, Math, and Underground collections of NCE, each of which contains three sequences of increasing difficulty and slightly different trajectories. First, we used [29] to build a prior model with the easy sequences and then ran our system in the medium and hard ones. Figure 1 shows the resulting prior model for the Math collection. We compared the performance of G-Loc with two SLAM (LIO-SAM [30], NV-LIO [38]) and two map-based localization methods (LiDAR-Loc [39], HDL-Loc [40]). For a fair comparison, all localization methods used the same prior map. Table III contains ATE RMSE, with the values for the SLAM methods extracted from [38]. The results showcase the robustness and accuracy of our system, which outperformed the localization methods and succeeded in every sequence. The failures of [39, 40] were caused by the lack of prior maps in certain regions (Quad-h) and aggressive motion (Math-h). G-Loc maintained accurate localization in these regions thanks to LIO, allowing for re-localization. G-Loc also outperformed the SLAM methods in most sequences, achieving an average error of 8 cm (note that the localization errors ultimately depend on the error of the prior model, see [29] for detailed values).
Method / Seq. Quad-m Quad-h Math-m Math-h Und-m Und-h Avg.
LIO-SAM [30] 0.067 0.13 0.13 0.085 0.065 0.42 0.15
NV-LIO [38] 0.076 0.18 0.095 0.094 0.072 0.078 0.10
LiDAR-Loc [39] 0.11 $\times$ 0.10 $\times$ 0.11 $\times$ 0.11
HDL-Loc [40] 0.10 $\times$ 0.11 $\times$ 0.11 0.12 0.11
G-Loc (ours) 0.067 0.087 0.090 0.093 0.074 0.075 0.081
TABLE III: ATE RMSE (m) in NCE dataset. Best results are bold and second best are underlined. $\times$ indicates an unrecoverable failure.
@@ V-B On the Time Consumption
While the above section aimed to demonstrate the robustness and precision of our approach, we now seek to demonstrate its real-time performance. The processing times per frame consumed by each algorithm in the conducted experiments are listed in Table II. G-Loc required, on average, $25$ ms in EULT and $34$ ms in CUZ —this is much faster than the typical LiDAR frame rate ($10$ Hz) and therefore meets real-time requirements. In both cases, LIO accounts for $60\%$ of that time, and the localization module for the rest (visualization and dynamic load are excluded as they operate in the background). G-Loc with NDTOmp took considerably longer, but the frame rate was generally not exceeded. We also report the time consumed by Autoware (marked with , as it used a better CPU), which is similar to G-Loc using NDTOmp.
Finally, we compared the use of G-Loc to running a full SLAM system to reveal the benefits of exploiting the known prior topo-metric map instead of building a new map. We used the EULT dataset because it is the largest and therefore where a new map would be more costly to generate and maintain. The time required to process a frame is similar to G-Loc (the LIO update), but the graph optimization and loop closing took $190.4\pm 101.4$ ms. These demanding processes slowed down the entire pipeline and caused interruptions in the localization.
@@ V-C Real-world Deployment
G-Loc has been deployed on an autonomous bus (see Figure 7a) on an urban bus line in the city of Zaragoza (Aragón, Spain) as part of the pioneering Spanish project on urban driving, DIGIZITY. The sensors installed on the vehicle and the route specifics appear in Table I. Our system requires coordination with other navigation systems to provide a robust and reliable localization, ensuring safe navigation even in areas with poor GNSS coverage. The prior model of the entire route is shown in Figure 7b. In addition, localization starts as soon as the bus is turned on. This means that the localization has to work without a prior model from the departure point until the route is reached ($2$ km). The system localizes in the unmapped region using LIO and GNSS in a sliding-graph optimization fashion until re-localization with the prior model occurs. Figure 7c depicts the trajectory estimated by G-Loc for a single trip, and Figure 7d shows the improvement of these estimates (blue) over the raw GNSS data (green) in one of the many low coverage areas. During the deployment, the bus operated seamlessly for about $100$ hours over a span of $20$ days, covering about $800$ km. Notably, there were no localization failures and the errors remained within bounds comparable to previous results (i.e., at the cm level), further showcasing the robustness and accuracy of G-Loc.
VI Conclusion
In this work we have presented G-Loc, a localization system that seamlessly integrates GNSS data, LIO and cloud-to-map registration into a graph optimization framework. We have proposed a novel method for map-based localization that leverages the results of a previous graph-SLAM process, utilizing both the geometric map and the underlying pose-graph representation. We have tested G-Loc in several datasets and in a real-world automotive application, demonstrating its accuracy, robustness, and adaptability to different sensor configurations and environmental conditions (e.g., the lack of prior knowledge in some parts of the environment). Long-term autonomy requires the sporadic execution of mapping to update the robot’s knowledge of the environment. This work sets the stage for developing a long-term approach where the prior model is updated and refined with online information.